last year asante drove his car 11,645 miles and changed his oil 4 times. if he drives 5,700 miles this year and changes his oil at the same rate, how many times will he change his oil this year?
Answer by rfer(16322)

4/11645=x/5700
11645x=22800
x=22800/11645
x=2 times